# Decremental $(1+\epsilon)$-Approximate Maximum Eigenvector: Dynamic Power Method

### v1

@@ 4 Conditional Lower Bounds for an Adaptive Adversary
Given an instance $(\delta,\kappa,\bm{\mathit{A}})$ of 1.3, we will run Algorithm 4 where Line 8 is implemented using $\mathcal{M}$. We will generate the input and the update sequence for $\mathcal{M}$ as follows. Set $\bm{\mathit{A}}_{0}\leftarrow\bm{\mathit{A}}$. Set $\epsilon$ and $T$ according to Algorithm 4. For $1\leq t\leq T$, we set $\bm{\mathit{v}}_{t}=\frac{1}{\sqrt{10}}\cdot\sqrt{\mu_{t-1}}\bm{\mathit{w}}_{t-1}$ according to Line 7 of Algorithm 4. We note that this is a valid update sequence for 1.1 since from Lemma 4.1, $\bm{\mathit{A}}_{t}=\bm{\mathit{A}}-\sum_{i=0}^{t-1}\frac{\mu_{i}}{10}\bm{\mathit{w}}_{i}\bm{\mathit{w}}_{i}^{\top}\succeq 0$.
Now, we describe what we return as an answer for 1.3. If at any time step $t$, we have $\mu_{t}<0$, then report that $\bm{\mathit{A}}$ is not psd. If Algorithm 4 returns False, then also report that $\bm{\mathit{A}}$ is not psd. Otherwise, we return $\bm{\mathit{X}}=\frac{1}{\sqrt{10}}\begin{bmatrix}\sqrt{\mu_{1}}\bm{\mathit{w}}_{1}&\sqrt{\mu_{2}}\bm{\mathit{w}}_{2}&\cdots&\sqrt{\mu_{T}}\bm{\mathit{w}}_{T}\end{bmatrix}$ as a certificate that $\bm{\mathit{A}}$ is psd. This completes the description of reduction from 1.3 to 1.1. The correctness follows directly from Lemma 4.2.
@@
5 Connections to Dynamic Positive Semi-definite Programs
We first define packing and covering semi-definite programs (SDPs). [^5] Some papers [Jam+21] flip our definition of packing and covering SDPs by considering their dual form.
Definition 5.1 (Packing/Covering SDP).
The covering SDP problem asks to find
Note that when the matrices $\bm{\mathit{C}}$ and $\bm{\mathit{A}}_{i}$’s are all diagonal matrices, the packing and covering SDP problems above are precisely the well-studied packing and covering LP problems, respectively. Near-linear time algorithms for $(1+\epsilon)$-approximately solving packing and covering LPs are very well-studied in a long line of work, some of which are [AO15, AO19, WRM16, Qua20], and these problems have many applications, such as in graph embedding [PST95], approximation algorithms [LN93, Tre98], scheduling [PST95], to name a few.
Dynamic LPs.
Near-optimal dynamic algorithms for packing and covering LPs were shown in [BKS23]. The paper studies two kinds of updates – restricting and relaxing updates. Restricting updates can only shrink the feasible region. In contrast, relaxing updates can only grow the feasible region. In [BKS23], the authors gave a deterministic algorithm that can maintain a $(1+\epsilon)$-approximate solution to either packing and covering LPs that undergo only restricting updates or only relaxing updates in total time $\widetilde{O}(N/\epsilon^{3}+t/\epsilon)$, where $N$ is the total number of nonzeros in the initial input and the updates, and $t$ is the number of updates. Hence, this is optimal up to logarithmic factors.
A natural question is whether one can generalize the near-optimal dynamic LP algorithms with polylogarithmic overhead by [BKS23] to work with SDPs since SDPs capture many further applications such as maximum cuts [IPS11, KL96], Sparse PCA [IPS11], sparsest cuts [IPS10], and balanced separators [OSV12], among many others.
Static SDPs.
Unfortunately, the algorithms for solving packing and covering SDPs are much more limited, even in the static setting. Near-linear time algorithms are known only for covering SDPs when the cost matrix $\bm{\mathit{C}}=\bm{\mathit{I}}$ is the identity [PTZ12, ALO16].
The fundamental barrier to working with general psd matrix $\bm{\mathit{C}}$ in covering SDPs is that it is as hard as approximating the minimum eigenvalue of $\bm{\mathit{C}}$ (consider the program $\max_{\bm{\mathit{Y}}\succeq 0}\operatorname{Tr}[\bm{\mathit{C}}\bm{\mathit{Y}}]$ such that $\operatorname{Tr}[\bm{\mathit{Y}}]\leq 1$). To the best of our knowledge, near-linear time algorithms for approximating the minimum eigenvalue assume a near-linear-time solver for $\bm{\mathit{C}}$, i.e., to compute $\bm{\mathit{C}}^{-1}\bm{\mathit{x}}$ in the near-linear time given $\bm{\mathit{x}}$. This can be done, for example, by applying the power method to $\bm{\mathit{C}}^{-1}$. When $\bm{\mathit{C}}^{-1}$ admits a fast solver, sometimes one can approximately solve a covering SDP fast, such as for spectral sparsification of graphs [LS17], and the max-cut problem [AK07, Tre09].
For packing SDPs, there is simply no near-linear time algorithm known. An algorithm for approximately solving packing SDPs and even the generalization to mixed packing-covering SDPs was claimed in [Jam+21], but there is an issue in the convergence analysis even for pure packing SDPs. Fast algorithms for this problem, hence, remain open.
Dynamic SDPs.
Problem 5.2 (Covering SDP with a Single Matrix Constraint under Restricting Updates).
Given $\bm{\mathit{A}}_{0}\succeq 0$, an accuracy parameter $\epsilon>0$, and an online sequence of vectors $\bm{\mathit{v}}_{1},\bm{\mathit{v}}_{2},\cdots,\bm{\mathit{v}}_{T}$ that update $\bm{\mathit{A}}_{t}\leftarrow\bm{\mathit{A}}_{t}-\bm{\mathit{v}}\bm{\mathit{v}}^{\top}$ such that $\bm{\mathit{A}}_{t}\succeq 0$. The problem asks to explicitly maintain, for all $t$, an $(1+\epsilon)$-approximate optimal value $\nu_{t}$ of the SDP, i.e.,
5.2 is equivalent to 1.1 in the following sense: given an algorithm for 1.1, we can obtain an algorithm for 5.2 with the same total update time and optimal query time. Conversely, given an algorithm for 5.2, we can obtain an algorithm for 1.1 in the eigenvalue-only version with the same total update time.
Proposition 5.3.
The following holds,
  1. 1.
    Given an algorithm for 1.1 with update time $\mathcal{T}$, there is an algorithm for Problem 5.2 with update time $\mathcal{T}$ and query time $O(n)$, i.e., the time required to query $\bm{\mathit{Q}}^{(t)}$ at any time $t$.
  2. 2.
    Given an algorithm for Problem 5.2 with update time $\widetilde{\mathcal{T}}$, there is an algorithm for 1.1 that only maintains the approximate eigenvalues with update time at most $\widetilde{\mathcal{T}}$.
Proof.
Let us first characterize the solution to Problem 5.2 at any $t$ for $\epsilon=0$. Since any feasible solution $\bm{\mathit{Y}}$ is PSD it can be characterized as, $\bm{\mathit{Y}}=\sum_{i=1}^{n}p_{i}\bm{\mathit{y}}_{i}\bm{\mathit{y}}_{i}^{\top}$ for some unit vectors $\bm{\mathit{y}}_{i}$’s. Now $\operatorname{Tr}[\bm{\mathit{Y}}]=\sum_{i}p_{i}$ and
We now show the first part, by proving that at any $t$, given $\epsilon$, the solution to 1.1 gives a solution to Problem 5.2. Let $\lambda_{t}$ and $\bm{\mathit{w}}_{t}$ denote an $\epsilon/2$-approximate solution to 1.1 for some $t$. Consider the solution $\bm{\mathit{Q}}^{(t)}=\frac{1}{\sqrt{(1-\epsilon/2)\lambda_{t}}}\bm{\mathit{w}}_{t}$ which gives $\bm{\mathit{Y}}^{(t)}=\frac{1}{(1-\epsilon/2)\lambda_{t}}\bm{\mathit{w}}_{t}\bm{\mathit{w}}_{t}^{\top}$. Then,
Therefore, $\bm{\mathit{Y}}^{(t)}$ satisfies the constraints of Problem 5.2. Next we look at the objective value, $\nu_{t}=\operatorname{Tr}[\bm{\mathit{Y}}^{(t)}]=\frac{1}{(1-\epsilon/2)\lambda_{t}}\leq\frac{1}{(1-\epsilon/2)^{2}\lambda_{\max}(\bm{\mathit{A}}_{t})}\leq(1+\epsilon)OPT_{t}.$ We can maintain $\bm{\mathit{Q}}^{(t)}$ by just maintaining $\bm{\mathit{w}}_{t},\lambda_{t}$ which requires no extra time. The time required to obtain $\bm{\mathit{Q}}^{(t)}$, which is the query time, from $\bm{\mathit{w}}_{t}$ and $\lambda_{t}$ is at most $O(nnz(\bm{\mathit{w}}_{t}))=O(n)$ and the value of $\nu_{t}$ can be obtained in $O(1)$ time.
To see the other direction, consider the solution $\bm{\mathit{Q}}^{(t)}$, $\nu_{t}$ of Problem 5.2. We can set $\lambda_{t}=\frac{1}{\nu_{t}}$. This implies that,
as required. In this case, we do not require any extra time. ∎
By plugging Theorem 1.2 into Proposition 5.3, we conclude the following.
Corollary 5.4.
There is a randomized algorithm for 5.2 under a sequence of $T$ restricting updates, that given $n,\bm{\mathit{A}}_{0}$ and $\epsilon>1/n$ as input, with probability at least $1-1/n$ works against an oblivious adversary in total update time
and query time $O(n)$.
@@ 6 Conclusion and Open Problems
As discussed in Section 5, Theorem 1.2 can be viewed as a starting point towards a dynamic algorithm for general positive semi-definite programs. Can we make further progress? The dynamic semi-definite program problem, even with just two matrix constraints, already seems to be challenging.
@@
One promising approach to attack this problem is to dynamize the matrix multiplicative weight update (MMWU) method for solving a packing/covering SDP [PTZ12] since the corresponding approach was successful for linear programs – the near-optimal algorithms of [BKS23] are essentially dynamized multiplicative weight update methods for positive linear programs. However, in our preliminary study exploring this approach, we could only obtain an algorithm that solves Problem 5.2, which has a single matrix constraint, and solves 1.1 partially, i.e., maintains an approximate eigenvalue only. The main barrier in this approach is that the algorithm requires maintaining the exponential of the sum of the constraint matrices, and to do this fast, we require that for any two constraint matrices $\bm{\mathit{A}}$ and $\bm{\mathit{B}}$, $e^{\bm{\mathit{A}}+\bm{\mathit{B}}}=e^{\bm{\mathit{A}}}e^{\bm{\mathit{B}}}$ which only holds when $\bm{\mathit{A}}$ and $\bm{\mathit{B}}$ commute i.e., $\bm{\mathit{A}}\bm{\mathit{B}}=\bm{\mathit{B}}\bm{\mathit{A}}$. Note that when $\bm{\mathit{A}}$ and $\bm{\mathit{B}}$ are diagonal, this is true; therefore, we can obtain the required algorithms for positive LPs. Even when we have just two constraint matrices where one of them is a diagonal matrix, this remains an issue as the matrices still do not commute.
Appendix A Omitted Proofs
@@
Algorithm 1 starts with a random vector $\bm{\mathit{v}}^{(0,r)}$ (Line 5). Let $l$ denote the dimension of the eigenspace corresponding to the maximum eigenvalue. We denote $\bm{\mathit{v}}^{(0,r)}=\sum_{i}\alpha^{(r)}_{i}\bm{\mathit{u}}_{i}$ and from Lemma A.1, we have that for all $r$, initial vectors,

### v2

@@ 4 Conditional Lower Bounds for an Adaptive Adversary
Given an instance $(\delta,\kappa,\bm{\mathit{A}})$ of 1.3, we will run Algorithm 4 where Line 8 is implemented using $\mathcal{M}$. We will generate the input and the update sequence for $\mathcal{M}$ as follows. Set $\bm{\mathit{A}}_{0}\leftarrow\bm{\mathit{A}}$. Set $\epsilon$ and $T$ according to Algorithm 4. For $1\leq t\leq T$, we set $\bm{\mathit{v}}_{t}=\frac{1}{\sqrt{10}}\cdot\sqrt{\mu_{t-1}}\bm{\mathit{w}}_{t-1}$ according to Line 7 of Algorithm 4. We note that this is a valid update sequence for 1.1 when $\bm{\mathit{A}}\succeq 0$ since from Lemma 4.1, if $\bm{\mathit{A}}\succeq 0$ then, $\bm{\mathit{A}}_{t}=\bm{\mathit{A}}-\sum_{i=0}^{t-1}\frac{\mu_{i}}{10}\bm{\mathit{w}}_{i}\bm{\mathit{w}}_{i}^{\top}\succeq 0$.
Now, we describe what we return as an answer for 1.3. From Lemma 4.2 if $\bm{\mathit{A}}$ is not PSD, then Algorithm 4 returns False and reports the matrix is not PSD. Additionally if $\bm{\mathit{A}}\succeq 0$, the algorithm returns $\bm{\mathit{X}}=\frac{1}{\sqrt{10}}\begin{bmatrix}\sqrt{\mu_{1}}\bm{\mathit{w}}_{1}&\sqrt{\mu_{2}}\bm{\mathit{w}}_{2}&\cdots&\sqrt{\mu_{T}}\bm{\mathit{w}}_{T}\end{bmatrix}$ as a certificate that $\bm{\mathit{A}}$ is PSD. This completes the reduction from 1.3 to 1.1.
@@ 5 Conclusion and Open Problems
Incremental Updates.
We believe that the corresponding incremental updates problem, i.e., we update the matrix as $\bm{\mathit{A}}_{t}\leftarrow\bm{\mathit{A}}_{t-1}+\bm{\mathit{v}}_{t}\bm{\mathit{v}}_{t}^{\top}$ cannot be solved in polylogarithmic amortized update time, even when the update sequence $\bm{\mathit{v}}_{t}$’s are from an oblivious adversary. At a high level, the incremental version of our problem seems significantly harder for the following reasons. When we perform decremental updates to a matrix, the new maximum eigenvector must be a part of the eigenspace spanned by the original maximum eigenvectors. Furthermore, it is easy to detect whether the maximum eigenvalue has gone down as we have shown in our paper. For the incremental setting, it is possible that after an update the maximum eigenvalue has gone up and the new maximum eigenvector is a direction that was not the previous one or the update direction and in such cases we cannot really detect this quickly with known information on previous eigenvectors and update directions. This can also happen $n$ times and in every such case, we have to compute the eigenvalue and eigenvector from scratch. Therefore, we leave lower bounds and algorithms for incremental setting as an open problem.
@@
As discussed in Appendix A, Theorem 1.2 can be viewed as a starting point towards a dynamic algorithm for general positive semi-definite programs. Can we make further progress? The dynamic semi-definite program problem, even with just two matrix constraints, already seems to be challenging.
@@
One promising approach to attack this problem is to dynamize the matrix multiplicative weight update (MMWU) method for solving a packing/covering SDP [peng2012faster] since the corresponding approach was successful for linear programs – the near-optimal algorithms of [bhattacharya2023dynamic] are essentially dynamized multiplicative weight update methods for positive linear programs. However, in our preliminary study exploring this approach, we could only obtain an algorithm that solves Problem A.2, which has a single matrix constraint, and solves 1.1 partially, i.e., maintains an approximate eigenvalue only. The main barrier in this approach is that the algorithm requires maintaining the exponential of the sum of the constraint matrices, and to do this fast, we require that for any two constraint matrices $\bm{\mathit{A}}$ and $\bm{\mathit{B}}$, $e^{\bm{\mathit{A}}+\bm{\mathit{B}}}=e^{\bm{\mathit{A}}}e^{\bm{\mathit{B}}}$ which only holds when $\bm{\mathit{A}}$ and $\bm{\mathit{B}}$ commute i.e., $\bm{\mathit{A}}\bm{\mathit{B}}=\bm{\mathit{B}}\bm{\mathit{A}}$. Note that when $\bm{\mathit{A}}$ and $\bm{\mathit{B}}$ are diagonal, this is true; therefore, we can obtain the required algorithms for positive LPs. Even when we have just two constraint matrices where one of them is a diagonal matrix, this remains an issue as the matrices still do not commute.
[…]
Appendix A Connections to Dynamic Positive Semi-definite Programs
We first define packing and covering semi-definite programs (SDPs). [^5] Some papers [jambulapati2021positive] flip our definition of packing and covering SDPs by considering their dual form.
Definition A.1 (Packing/Covering SDP).
The covering SDP problem asks to find
Note that when the matrices $\bm{\mathit{C}}$ and $\bm{\mathit{A}}_{i}$’s are all diagonal matrices, the packing and covering SDP problems above are precisely the well-studied packing and covering LP problems, respectively. Near-linear time algorithms for $(1+\epsilon)$-approximately solving packing and covering LPs are very well-studied in a long line of work, some of which are [allen2015nearly, allen2019nearly, wang2016unified, quanrud2020nearly], and these problems have many applications, such as in graph embedding [plotkin1995fast], approximation algorithms [luby1993parallel, trevisan1998parallel], scheduling [plotkin1995fast], to name a few.
Dynamic LPs.
Near-optimal dynamic algorithms for packing and covering LPs were shown in [bhattacharya2023dynamic]. The paper studies two kinds of updates – restricting and relaxing updates. Restricting updates can only shrink the feasible region. In contrast, relaxing updates can only grow the feasible region. In [bhattacharya2023dynamic], the authors gave a deterministic algorithm that can maintain a $(1+\epsilon)$-approximate solution to either packing and covering LPs that undergo only restricting updates or only relaxing updates in total time $\widetilde{O}(N/\epsilon^{3}+t/\epsilon)$, where $N$ is the total number of nonzeros in the initial input and the updates, and $t$ is the number of updates. Hence, this is optimal up to logarithmic factors.
A natural question is whether one can generalize the near-optimal dynamic LP algorithms with polylogarithmic overhead by [bhattacharya2023dynamic] to work with SDPs since SDPs capture many further applications such as maximum cuts [iyengar2011approximating, klein1996efficient], Sparse PCA [iyengar2011approximating], sparsest cuts [iyengar2010feasible], and balanced separators [orecchia2012approximating], among many others.
Static SDPs.
Unfortunately, the algorithms for solving packing and covering SDPs are much more limited, even in the static setting. Near-linear time algorithms are known only for covering SDPs when the cost matrix $\bm{\mathit{C}}=\bm{\mathit{I}}$ is the identity [peng2012faster, allen2016using].
The fundamental barrier to working with general psd matrix $\bm{\mathit{C}}$ in covering SDPs is that it is as hard as approximating the minimum eigenvalue of $\bm{\mathit{C}}$ (consider the program $\max_{\bm{\mathit{Y}}\succeq 0}\operatorname{Tr}[\bm{\mathit{C}}\bm{\mathit{Y}}]$ such that $\operatorname{Tr}[\bm{\mathit{Y}}]\leq 1$). To the best of our knowledge, near-linear time algorithms for approximating the minimum eigenvalue assume a near-linear-time solver for $\bm{\mathit{C}}$, i.e., to compute $\bm{\mathit{C}}^{-1}\bm{\mathit{x}}$ in the near-linear time given $\bm{\mathit{x}}$. This can be done, for example, by applying the power method to $\bm{\mathit{C}}^{-1}$. When $\bm{\mathit{C}}^{-1}$ admits a fast solver, sometimes one can approximately solve a covering SDP fast, such as for spectral sparsification of graphs [lee2017sdp], and the max-cut problem [arora2007combinatorial, trevisan2009max].
For packing SDPs, there is simply no near-linear time algorithm known. An algorithm for approximately solving packing SDPs and even the generalization to mixed packing-covering SDPs was claimed in [jambulapati2021positive], but there is an issue in the convergence analysis even for pure packing SDPs. Fast algorithms for this problem, hence, remain open.
Dynamic SDPs.
Problem A.2 (Covering SDP with a Single Matrix Constraint under Restricting Updates).
Given $\bm{\mathit{A}}_{0}\succeq 0$, an accuracy parameter $\epsilon>0$, and an online sequence of vectors $\bm{\mathit{v}}_{1},\bm{\mathit{v}}_{2},\cdots,\bm{\mathit{v}}_{T}$ that update $\bm{\mathit{A}}_{t}\leftarrow\bm{\mathit{A}}_{t}-\bm{\mathit{v}}_{t}\bm{\mathit{v}}_{t}^{\top}$ such that $\bm{\mathit{A}}_{t}\succeq 0$. The problem asks to explicitly maintain, for all $t$, an $(1+\epsilon)$-approximate optimal value $\nu_{t}$ of the SDP, i.e.,
A.2 is equivalent to 1.1 in the following sense: given an algorithm for 1.1, we can obtain an algorithm for A.2 with the same total update time and optimal query time. Conversely, given an algorithm for A.2, we can obtain an algorithm for 1.1 in the eigenvalue-only version with the same total update time.
Proposition A.3.
The following holds,
  1. 1.
    Given an algorithm for 1.1 with update time $\mathcal{T}$, there is an algorithm for Problem A.2 with update time $\mathcal{T}$ and query time $O(n)$, i.e., the time required to query $\bm{\mathit{Q}}^{(t)}$ at any time $t$.
  2. 2.
    Given an algorithm for Problem A.2 with update time $\widetilde{\mathcal{T}}$, there is an algorithm for 1.1 that only maintains the approximate eigenvalues with update time at most $\widetilde{\mathcal{T}}$.
Proof.
Let us first characterize the solution to Problem A.2 at any $t$ for $\epsilon=0$. Since any feasible solution $\bm{\mathit{Y}}$ is PSD it can be characterized as, $\bm{\mathit{Y}}=\sum_{i=1}^{n}p_{i}\bm{\mathit{y}}_{i}\bm{\mathit{y}}_{i}^{\top}$ for some unit vectors $\bm{\mathit{y}}_{i}$’s. Now $\operatorname{Tr}[\bm{\mathit{Y}}]=\sum_{i}p_{i}$ and
We now show the first part, by proving that at any $t$, given $\epsilon$, the solution to 1.1 gives a solution to Problem A.2. Let $\lambda_{t}$ and $\bm{\mathit{w}}_{t}$ denote an $\epsilon/2$-approximate solution to 1.1 for some $t$. Consider the solution $\bm{\mathit{Q}}^{(t)}=\frac{1}{\sqrt{(1-\epsilon/2)\lambda_{t}}}\bm{\mathit{w}}_{t}$ which gives $\bm{\mathit{Y}}^{(t)}=\frac{1}{(1-\epsilon/2)\lambda_{t}}\bm{\mathit{w}}_{t}\bm{\mathit{w}}_{t}^{\top}$. Then,
Therefore, $\bm{\mathit{Y}}^{(t)}$ satisfies the constraints of Problem A.2. Next we look at the objective value, $\nu_{t}=\operatorname{Tr}[\bm{\mathit{Y}}^{(t)}]=\frac{1}{(1-\epsilon/2)\lambda_{t}}\leq\frac{1}{(1-\epsilon/2)^{2}\lambda_{\max}(\bm{\mathit{A}}_{t})}\leq(1+\epsilon)OPT_{t}.$ We can maintain $\bm{\mathit{Q}}^{(t)}$ by just maintaining $\bm{\mathit{w}}_{t},\lambda_{t}$ which requires no extra time. The time required to obtain $\bm{\mathit{Q}}^{(t)}$, which is the query time, from $\bm{\mathit{w}}_{t}$ and $\lambda_{t}$ is at most $O(nnz(\bm{\mathit{w}}_{t}))=O(n)$ and the value of $\nu_{t}$ can be obtained in $O(1)$ time.
To see the other direction, consider the solution $\bm{\mathit{Q}}^{(t)}$, $\nu_{t}$ of Problem A.2. We can set $\lambda_{t}=\frac{1}{\nu_{t}}$. This implies that,
as required. In this case, we do not require any extra time. ∎
By plugging Theorem 1.2 into Proposition A.3, we conclude the following.
Corollary A.4.
There is a randomized algorithm for A.2 under a sequence of $T$ restricting updates, that given $n,\bm{\mathit{A}}_{0}$ and $\epsilon>1/n$ as input, with probability at least $1-1/n$ works against an oblivious adversary in total update time
and query time $O(n)$.
Appendix B Omitted Proofs
@@
Algorithm 1 starts with a random vector $\bm{\mathit{v}}^{(0,r)}$ (Line 5). Let $l$ denote the dimension of the eigenspace corresponding to the maximum eigenvalue. We denote $\bm{\mathit{v}}^{(0,r)}=\sum_{i}\alpha^{(r)}_{i}\bm{\mathit{u}}_{i}$ and from Lemma B.1, we have that for all $r$, initial vectors,